# Humanoid Occupancy: Enabling A Generalized Multimodal Occupancy Perception System on Humanoid Robots

### v1

@@ 4.2 Main results
FBOcc (Li et al., 2023b) C 1 76.8M 47.37 59.1 39.78 55.43 42.95 45.37 62.69 51.19 35.73 44.57 59.12 49.36 53.99 28.27
BEVFusion (Liu et al., 2022) C+L 1 60.6M 53.98 60.24 51 51.7 36.72 43.31 66.64 63.38 41.42 71.4 65.29 61.38 62.96 32.45
HumaniodOcc (Ours) C+L 1 40.5M 52.79 60.49 49.23 49.22 35.28 43.63 57.66 63.9 43.07 71.41 64.17 60.89 59.43 35.54
BEVDet (Huang et al., 2021) C 2 75.3M 47.95 55 39.37 55.66 28.32 39.29 62.42 51.56 36.85 58.61 63.85 54.96 55.03 29.52
FBOcc (Li et al., 2023b) C 2 76.8M 46.7 58.45 40.11 53.52 42.97 45.44 62.14 50.83 35.38 41.06 58.08 49.07 54 27.76
BEVFusion (Liu et al., 2022) C+L 2 60.6M 53.21 59.56 51 52.16 37.03 40.83 57.63 63.88 44.38 71.68 63.87 62.45 61.16 32.43
HumaniodOcc (Ours) C+L 2 40.5M 55.73 61.32 50.85 56.16 40.49 45.49 65.6 64.76 44.07 72.77 65.58 63.01 64.5 35.48

### v2

@@ 4.2 Main results
FBOcc (Li et al., 2023b) C 1 76.8M 47.37 59.10 39.78 55.43 42.95 45.37 62.69 51.19 35.73 44.57 59.12 49.36 53.99 28.27
BEVFusion (Liu et al., 2022) C+L 1 60.6M 53.98 60.24 51.00 51.70 36.72 43.31 66.64 63.38 41.42 71.40 65.29 61.38 62.96 32.45
HumaniodOcc (Ours) C+L 1 40.5M 52.79 60.49 49.23 49.22 35.28 43.63 57.66 63.90 43.07 71.41 64.17 60.89 59.43 35.54
BEVDet (Huang et al., 2021) C 2 75.3M 47.95 55.00 39.37 55.66 28.32 39.29 62.42 51.56 36.85 58.61 63.85 54.96 55.03 29.52
FBOcc (Li et al., 2023b) C 2 76.8M 46.70 58.45 40.11 53.52 42.97 45.44 62.14 50.83 35.38 41.06 58.08 49.07 54.00 27.76
BEVFusion (Liu et al., 2022) C+L 2 60.6M 53.21 59.56 51.00 52.16 37.03 40.83 57.63 63.88 44.38 71.68 63.87 62.45 61.16 32.43
HumaniodOcc (Ours) C+L 2 40.5M 55.73 61.32 50.85 56.16 40.49 45.49 65.60 64.76 44.07 72.77 65.58 63.01 64.50 35.48